# subalign[1][]

### v1

@@ Abstract
Nuclei far from the valley of stability provide a valuable laboratory for probing the dependence on nuclear forces and the nature of the quantum correlations among nucleons. There have been several ab initio calculations of neutron-rich oxygen isotopes Hagen et al. (2009); Otsuka et al. (2010); Hagen et al. (2012); Holt et al. (2013); Cipollone et al. (2013); Jansen et al. (2014); Stroberg et al. (2021); Ma et al. (2020); Kaur et al. (2022) as well as neutron-rich carbon isotopes Forssen et al. (2011); Jansen et al. (2014); Kanungo et al. (2016); Tran et al. (2018); Stroberg et al. (2021); Li et al. (2024). In this work, we perform calculations of neutron-rich carbon and oxygen isotopes using nuclear lattice effective field theory (NLEFT). We use chiral effective field theory (EFT) interactions defined on a three-dimensional lattice and perform quantum Monte Carlo simulations of the many-body system using auxiliary fields. Reviews of NLEFT and related methods can be found in Refs. Lee (2009); Drut and Nicholson (2013); Lee (2017); Lähde and Meißner (2019), and reviews of chiral EFT can be found in Refs. Epelbaum et al. (2009); Machleidt and Sammarruca (2016); Hammer et al. (2020).
@@
Wavefunction matching was introduced in Ref. Elhatisari et al. (2024) to accelerate the convergence of perturbation theory. We also use wavefunction matching in this work and apply the interactions defined in Ref. Elhatisari et al. (2024) with spatial lattice spacing $a=1.32$ fm. Details of the interactions and computational methods can be found in the Supplemental Material accompanying Ref. Elhatisari et al. (2024). For our chiral interactions, a low-energy scheme is used where the two-nucleon two-pion exchange and higher-pion exchange interactions are treated as short-range contact interactions. Within this framework, we include all two-nucleon and three-nucleon interactions up to $O(Q^{4})$ or next-to-next-to-next-to-leading order (N3LO). This includes chiral three-nucleon interactions such as the one-pion exchange, two-pion exchange, and short-range three-nucleon interactions. As introduced in Ref. Elhatisari et al. (2024), we also include additional three-nucleon interactions that correspond with specific choices for the local regulators used in the three-nucleon interactions. We have not included any four-nucleon interactions.
@@
In Fig. 1, we present lattice results for the energies of the neutron-rich carbon and oxygen isotopes versus the number of nucleons, $A$. The energies for 12-14C, of the first two excited states in 12C and 16-18O were already reported in Ref. Elhatisari et al. (2024), and they are shown again in the results here. The error bars correspond to one standard deviation and include statistical errors as well as uncertainties in the extrapolation to infinite Euclidean time and infinite volume. While there are some small deviations in comparison with experimental data, the overall agreement is quite good. In future work, we plan to investigate the remaining sources of errors and perform calculations of other observables such as charge radii, quadrupole moments, electromagnetic transitions, and magnetic dipole moments.
@@
Figure 1: Ground state energies for the neutron-rich carbon and oxygen isotopes. NLEFT results at order N3LO are compared with experimental data. In the case of 12C, we also show the first two excited states.
@@
Having demonstrated that the lattice calculations accurately reproduce the energies of the neutron-rich carbon and oxygen isotopes, we now turn our attention to probing the dependence on nuclear forces and measuring quantum correlations. In each partial-wave channel, we calculate $\braket{\Psi}{\Delta O}{\Psi}$ for some perturbing two-nucleon operator $\Delta O$. Similar sensitivity studies have been performed in the literature Ekström and Hagen (2019); Hu et al. (2022). In our analysis, however, we do not focus on the details of $\Delta O$ but rather the change to the scattering phase shifts, $\Delta\delta(p)$. By relying on physical observables, we are constructing a model-independent framework that can be translated to any low-energy EFT calculation. Two different EFT calculations would simply agree on $\Delta\delta(p)$ and determine their corresponding operators $\Delta O$ accordingly. Induced higher-body operators can also be determined by matching to higher-body physical observables.
@@
For each partial-wave channel, we consider a short-range two-nucleon interaction operator that, when added to the full Hamiltonian, produces a $1\%$ reduction in the scattering phase shift at relative momentum $p=150$ MeV. The detailed form of the operators we use and their effect on the scattering phase shifts are described in the Supplemental Material SM . Before presenting lattice results for the two-nucleon correlations, we first prove a useful fact about isospin correlations that we call $T_{z}$ linearity.
Let |$\ket{\Psi_{(1/2,-1/2)}}$ be a nuclear state with isospin $T=1/2$ and $T_{z}=-1/2$. For example, |$\ket{\Psi_{(1/2,-1/2)}}$ could be the ground of a nucleus such as 13C or 17O with one more neutron than the number of protons. Let $A_{(1,T_{z})}$ be an operator with isospin $T=1$ and arbitrary $T_{z}$. For example, $A_{(1,T_{z})}$ could be a short-range operator that annihilates two nucleons in some $T=1$ partial-wave channel. Then $T_{z}=-1$ corresponds to the annihilation of two protons, $T_{z}=0$ corresponds to the isospin-symmetric annihilation of a proton and neutron, and $T_{z}=1$ corresponds to the annihilation of two neutrons. We now consider the operator expectation value,
$f(T_{z})=\braket{\Psi_{(1/2,-1/2)}}{A^{\dagger}_{(1,T_{z})}A_{(1,T_{z})}}{\Psi_{(1/2,-1/2)}}.$ (3)
We note that $A_{(1,T_{z})}\ket{\Psi_{(1/2,-1/2)}}$ can be decomposed into two irreducible isospin representations, $T=3/2$ and $T=1/2$. Let us write $f_{3/2}$ for the $3/2$ amplitude and $f_{1/2}$ for the $1/2$ amplitude. It is straightforward to show that $f(-1)=f_{3/2}$, $f(0)=\frac{2}{3}f_{3/2}+\frac{1}{3}f_{1/2}$, and $f(1)=\frac{1}{3}f_{3/2}+\frac{2}{3}f_{1/2}$. Therefore, the dependence on $T_{z}$ is linear, and we have the relation $f(1)=2f(0)-f(-1)$.
@@
We now compare $f(1)$ against the linear combination $2f(0)-f(-1)$. If each of the extra neutrons are uncorrelated with each other, then the additional correlations produced by each extra neutron are additive, and we expect $T_{z}$ linearity to still hold, $f(1)=2f(0)-f(-1)$. In general, however, there will be some correlations among the extra neutrons, and this results in $f(1)$ being different from $2f(0)-f(-1)$. The comparison between $f(1)$ and $2f(0)-f(-1)$ is therefore a measure of correlations among the extra neutrons in a neutron-rich nucleus.
@@
In Fig. 2, we show 1S0 correlations for the combinations proton-proton (pp), proton-neutron (pn), neutron-neutron (nn), and twice proton-neutron minus proton-proton (2pn--pp). The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes. In both cases, the pp correlations are rather flat, decreasing by only $14\%$ from 16O to 26O and decreasing only $15\%$ from 12C to 23C. This is an indication that the proton structure of the nuclear core does not change much. Previous lattice simulations have shown that the ground states of 16O and 12C both have significant alpha cluster substructures Epelbaum et al. (2011, 2012, 2014); Freer et al. (2018); Summerfield et al. (2021); Shen et al. (2021, 2023). Our results here suggest that the pp correlations within the alpha clusters remain mostly intact as extra neutrons are added.
@@
We see that the 1S0 nn correlations for oxygen and carbon both have a prominent “staircase” pattern produced by superfluid pairing. We note that the pp, pn, nn correlations are equal for 16O and for 12C due to isospin symmetry. Due to $T_{z}$ linearity, we observe that the nn correlations equal the 2pn--pp correlations for 17O and for 13C. In each of the correlation measurements presented here, we have not included perturbative theory corrections to the correlations. Therefore, the correlations being measured are those associated with the nonperturbative Hamiltonian used in the propagation of the wavefunction, and the nonperturbative Hamiltonian used has exact isospin symmetry.
@@
If we look closely at the 1S0 nn correlations for oxygen and carbon, we see that adding an unpaired or odd neutron produces an increase in $\Delta E$ whose slope matches that of 2pn--pp. See, for example, the increase from 18O to 19O, 20O to 21O, 14C to 15C, or 16C to 17C. A simple interpretation of this result is that the unpaired neutron is only weakly correlated with the other extra neutrons and is predominantly interacting with the $T=0$ nuclear core. On the other hand, adding one more neutron to complete the 1S0 pair produces an increase in $\Delta E$ with slope rising higher than that of 2pn--pp. This additional neutron is interacting strongly with its pair partner as well as with the nuclear core. We note that the pn correlations follow a smooth and almost linear trajectory as a function of the number of neutrons.
@@
Figure 2: Correlations for pp, pn, nn, and 2pn--pp in the 1S0 channel. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes.
@@
In Fig. 3, we show 3P0 correlations for pp, pn, nn, and 2pn--pp. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes. We again note that the pp, pn, nn correlations are equal for 16O and 12C due to isospin symmetry, and the nn and 2pn--pp correlations are equal for 17O and 13C due to $T_{z}$ linearity. We observe that the 3P0 pp correlations decrease gradually with the number of neutrons, but at a faster rate than we observed for the 1S0 channel. The decrease is $25\%$ from 16O to 26O, and the decrease is $49\%$ from 12C to 23C. We note that P-wave correlations between protons would not come from protons within one alpha cluster, but rather protons from two different neighboring alpha clusters. These results suggest that while the alpha clusters may remain intact, they may become less correlated with each other as extra neutrons are added.
@@
The nn correlations for the oxygen isotopes remain very close to the 2pn--pp correlations even for up to six extra neutrons. The same is true for the carbon isotopes for up to four extra neutrons. The 3P0 correlations between the extra neutrons grow slowly and smoothly with the number of neutrons. The same is true for the other $T=1$ partial waves. We note that there are some faint oscillations in the P-wave correlations due to the pairing driven by the 1S0 interactions. In the Supplemental Material SM , we present results for the other partial waves, including both $T=1$ and $T=0$ channels.
@@
There has been considerable discussion in the recent literature about short-range correlations and $T=0$ proton-neutron pairs Hen et al. (2014); Duer et al. (2018); Schmookler et al. (2019); Cruz-Torres et al. (2021); Tropiano et al. (2021). These short-range correlations arise from the singular tensor force and depend strongly on the short-distance resolution scale. In our calculations, we have used a relatively low resolution scale associated with our $1.32$ fm lattice spacing, and the total $T=0$ S-wave correlations are larger than the total $T=1$ S-wave correlations by only $26\%$ for 16O and only $25\%$ for 12C. The near equality of the $T=0$ and $T=1$ contributions is related to the hidden spin-isospin exchange symmetry discussed in Ref. Lee et al. (2021).
@@
Figure 3: Correlations for pp, pn, nn, and 2pn--pp in the 3P0 channel. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes.
@@
We have presented ab initio lattice results for the neutron-rich carbon and oxygen isotopes using high-fidelity chiral interactions. The energies are in good agreement with experimental data. We have also computed correlations associated with two-nucleon interaction operators in various partial-wave channels. By studying the dependence on $T_{z}$ in the $T=1$ channels, we are able to measure correlations among the extra neutrons in the neutron-rich carbon and oxygen isotopes. For the 1S0 channel, we find that any paired neutron interacts with the nuclear core and its neutron pair partner, while any unpaired neutron interacts primarily with only the nuclear core. For the other partial waves, the correlations among the extra neutrons grow slowly and smoothly with the number of neutrons. These findings for the carbon and oxygen isotopes may in fact be universal properties that can be seen many other neutron-rich nuclei.
@@
The observed “staircase” pattern for the 1S0 nn correlations may have an impact on the charge radii for the carbon and oxygen isotopes with even and odd numbers of neutrons. We plan to investigate these effects in the future using the pinhole algorithm Elhatisari et al. (2017). However, pinhole calculations of $A$-body density correlations do not have an immediate analog for other nuclear many-body methods. It is therefore valuable that a significant amount of information about nuclear forces and quantum correlations can be deduced from the simple correlation measurements presented here and can be expressed in a model-independent language. The correlation studies presented here can be readily adopted by other groups using other nuclear many-body methods.

### v2

# Ab Initio Calculations of the Carbon and Oxygen Isotopes: Energies, Correlations, and Superfluid Pairing

@@ Abstract
keywords:
Nuclear Lattice Effective Field Theory, Neutron-rich Isotopes , Superfluid parings , Ab Initio calculation
journal: Physics Letters B
@@
Nuclei far from the valley of stability provide a valuable laboratory for probing the dependence on nuclear forces and the nature of the quantum correlations among nucleons. There have been several ab initio calculations of neutron-rich oxygen isotopes [Hagen:2009mm, Otsuka:2009cs, Hagen:2012sh, Holt:2012fr, Cipollone:2013zma, Jansen:2014qxa, Stroberg:2019bch, Ma:2020spz, Kaur:2022yoh] as well as neutron-rich carbon isotopes [Forssen:2011dr, Jansen:2014qxa, Kanungo:2016tmz, Tran:2017gxv, Stroberg:2019bch, Li:2024kno]. In this work, we perform calculations of neutron-rich carbon and oxygen isotopes using nuclear lattice effective field theory (NLEFT). We use chiral effective field theory (EFT) interactions defined on a three-dimensional lattice and perform quantum Monte Carlo simulations of the many-body system using auxiliary fields. Reviews of NLEFT and related methods can be found in Refs. [Lee:2008fa, Drut:2012md, Lee:2016fhn, Lahde:2019npb], and reviews of chiral EFT can be found in Refs. [Epelbaum:2008ga, Machleidt:2016rvv, Hammer:2019poc].
@@
Wavefunction matching was introduced in Ref. [Elhatisari:2022zrb] to accelerate the convergence of perturbation theory. We also use wavefunction matching in this work and apply the interactions defined in Ref. [Elhatisari:2022zrb] with spatial lattice spacing $a=1.32$ fm. Details of the interactions and computational methods can be found in the Supplemental Material accompanying Ref. [Elhatisari:2022zrb]. For our chiral interactions, a low-energy scheme is used where the two-nucleon two-pion exchange and higher-pion exchange interactions are treated as short-range contact interactions. Within this framework, we include all two-nucleon and three-nucleon interactions up to $O(Q^{4})$ or next-to-next-to-next-to-leading order (N3LO). This includes chiral three-nucleon interactions such as the one-pion exchange, two-pion exchange, and short-range three-nucleon interactions. As introduced in Ref. [Elhatisari:2022zrb], we also include additional three-nucleon interactions that correspond with specific choices for the local regulators used in the three-nucleon interactions. We have not included any four-nucleon interactions.
@@
In Fig. 1, we present lattice results for the energies of the neutron-rich carbon and oxygen isotopes versus the number of nucleons, $A$. The energies for 12-14C, of the first two excited states in 12C and 16-18O were already reported in Ref. [Elhatisari:2022zrb], and they are shown again in the results here. The error bars correspond to one standard deviation and include statistical errors as well as uncertainties in the extrapolation to infinite Euclidean time and infinite volume. While there are some small deviations in comparison with experimental data, the overall agreement is quite good. In future work, we plan to investigate the remaining sources of errors and perform calculations of other observables such as charge radii, quadrupole moments, electromagnetic transitions, and magnetic dipole moments.
@@
Figure 1: Ground state energies for the neutron-rich carbon and oxygen isotopes. NLEFT results at order N3LO are compared with experimental data. In the case of 12C, we also show the first two excited states. Experimental data are from [Kelley:2017qgh] [Wang:2021xhn].
@@
Having demonstrated that the lattice calculations accurately reproduce the energies of the neutron-rich carbon and oxygen isotopes, we now turn our attention to probing the dependence on nuclear forces and measuring quantum correlations. In each partial-wave channel, we calculate $\braket{\Psi|\Delta O|\Psi}$ for some perturbing two-nucleon operator $\Delta O$. Similar sensitivity studies have been performed in the literature [Ekstrom:2019lss, Hu:2021trw]. In our analysis, however, we do not focus on the details of $\Delta O$ but rather the change to the scattering phase shifts, $\Delta\delta(p)$. By relying on physical observables, we are constructing a model-independent framework that can be translated to any low-energy EFT calculation. We mean ‘model-independent’ in the sense that one can compare two EFTs regardless of their details in operator structure by matching low-energy observables. Two different EFT calculations would simply agree on $\Delta\delta(p)$ and determine their corresponding operators $\Delta O$ accordingly. Induced higher-body operators can also be determined by matching to higher-body physical observables.
@@
For each partial-wave channel, we consider a short-range two-nucleon interaction operator that, when added to the full Hamiltonian, produces a $1\%$ reduction in the scattering phase shift at relative momentum $p=150$ MeV. The detailed form of the operators we use and their effect on the scattering phase shifts are described in the Supplemental Material [SM]. Before presenting lattice results for the two-nucleon correlations, we first prove a useful fact about isospin correlations that we call $T_{z}$ linearity. More specifically, we will consider the ‘energy correlation’ between two nucleons.
[…]
$f(T_{z})=\braket{\Psi_{(1/2,-1/2)}|A^{\dagger}_{(1,T_{z})}A_{(1,T_{z})}|\Psi_{(1/2,-1/2)}}.$ (1)
[…]
@@
We now compare $f(1)$ against the linear combination 2f(0)f(1)2f(0)-f(-1). If each of the extra neutrons are uncorrelated with each other, then the additional correlations produced by each extra neutron are additive, and we expect $T_{z}$ linearity to still hold, f(1)=2f(0)f(1)f(1)=2f(0)-f(-1). In general, however, there will be some correlations among the extra neutrons, and this results in $f(1)$ being different from 2f(0)f(1)2f(0)-f(-1). The comparison between $f(1)$ and 2f(0)f(1)2f(0)-f(-1) is therefore a measure of correlations among the extra neutrons in a neutron-rich nucleus.
@@
In Fig. 2, we show 1S0 correlations for the combinations proton-proton (pp), proton-neutron (pn), neutron-neutron (nn), and twice proton-neutron minus proton-proton (2pn-$p$). The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes. In both cases, the pp correlations are rather flat, decreasing by only $14\%$ from 16O to 26O and decreasing only $15\%$ from 12C to 23C. This is an indication that the proton structure of the nuclear core does not change much. Previous lattice simulations have shown that the ground states of 16O and 12C both have significant alpha cluster substructures [Epelbaum:2011md, Epelbaum:2012qn, Epelbaum:2013paa, Freer:2017gip, Summerfield:2021oex, Shen:2021kqr, Shen:2022bak]. Our results here suggest that the pp correlations within the alpha clusters remain mostly intact as extra neutrons are added.
@@
We see that the 1S0 nn correlations for oxygen and carbon both have a prominent “staircase” pattern produced by superfluid pairing. We note that the pp, pn, nn correlations are equal for 16O and for 12C due to isospin symmetry. Due to $T_{z}$ linearity, we observe that the nn correlations equal the 2pn-$p$ correlations for 17O and for 13C. In each of the correlation measurements presented here, we have not included perturbative theory corrections in nonperturbative wave function to the correlations. Therefore, the correlations being measured are those associated with the nonperturbative Hamiltonian used in the propagation of the wavefunction, and the nonperturbative Hamiltonian used has exact isospin symmetry.
@@
If we look closely at the 1S0 nn correlations for oxygen and carbon, we see that adding an unpaired or odd neutron produces an increase in $\Delta E$ whose slope is very close to that of 2pn-pp. See, for example, the increase from 18O to 19O, 20O to 21O, 14C to 15C, or 16C to 17C. A simple interpretation of this result is that the unpaired neutron is only weakly correlated with the other extra neutrons and is predominantly interacting with the $T=0$ nuclear core. On the other hand, adding one more neutron to complete the 1S0 pair produces an increase in $\Delta E$ with slope rising higher than that of 2pn-pp. This additional neutron is interacting strongly with its pair partner as well as with the nuclear core. We note that the pn correlations follow a smooth and almost linear trajectory as a function of the number of neutrons. See Ref. [Brown:2013] for a discussion of pairing for these nuclei in the shell model.
@@
Figure 2: Correlations for pp, pn, nn, and 2pn-$p$ in the 1S0 channel. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes.
@@
In Fig. 3, we show 3P0 correlations for pp, pn, nn, and 2pn-pp. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes. We again note that the pp, pn, nn correlations are equal for 16O and 12C due to isospin symmetry, and the nn and 2pn-$p$ correlations are equal for 17O and 13C due to $T_{z}$ linearity. We observe that the 3P0 pp correlations decrease gradually with the number of neutrons, but at a faster rate than we observed for the 1S0 channel. The decrease is $25\%$ from 16O to 26O, and the decrease is $49\%$ from 12C to 23C. We note that P-wave correlations between protons would not come from protons within one alpha cluster, but rather protons from two different neighboring alpha clusters. These results suggest that while the alpha clusters may remain intact, they may become less correlated with each other as extra neutrons are added.
@@
The nn correlations for the oxygen isotopes remain very close to the 2pn-$p$ correlations even for up to six extra neutrons. The same is true for the carbon isotopes for up to four extra neutrons. The 3P0 correlations between the extra neutrons grow slowly and smoothly with the number of neutrons. The same is true for the other $T=1$ partial waves. We note that there are some faint oscillations in the P-wave correlations due to the pairing driven by the 1S0 interactions. In the Supplemental Material [SM], we present results for the other partial waves, including both $T=1$ and $T=0$ channels.
@@
There has been considerable discussion in the recent literature about short-range correlations and $T=0$ proton-neutron pairs [Hen:2014nza, CLAS:2018yvt, CLAS:2019vsb, Cruz-Torres:2019fum, Tropiano:2021qgf]. These short-range correlations arise from the singular tensor force and depend strongly on the short-distance resolution scale. In our calculations, we have used a relatively low resolution scale associated with our $1.32$ fm lattice spacing, and the total $T=0$ S-wave correlations are larger than the total $T=1$ S-wave correlations by only $26\%$ for 16O and only $25\%$ for 12C. The near equality of the $T=0$ and $T=1$ contributions is related to the hidden spin-isospin exchange symmetry discussed in Ref. [Lee:2020esp].
@@
Figure 3: Correlations for pp, pn, nn, and 2pn-$p$ in the 3P0 channel. The top panel shows the oxygen isotopes, and the bottom panel shows the carbon isotopes.
@@
We have presented ab initio lattice results for the neutron-rich carbon and oxygen isotopes using high-fidelity chiral interactions. The energies are in good agreement with experimental data. We have also computed correlations associated with two-nucleon interaction operators in various partial-wave channels. By studying the dependence on $T_{z}$ in the $T=1$ channels, we are able to measure correlations among the extra neutrons in the neutron-rich carbon and oxygen isotopes. For the 1S0 channel, we find that any paired neutron interacts with the nuclear core and its neutron pair partner, while any unpaired neutron interacts primarily with only the nuclear core. For the other partial waves, the correlations among the extra neutrons grow slowly and smoothly with the number of neutrons. These findings for the carbon and oxygen isotopes may in fact be universal properties that can be seen in many other neutron-rich nuclei.
@@
The observed “staircase” pattern for the 1S0 nn correlations may have an impact on the charge radii for the carbon and oxygen isotopes with even and odd numbers of neutrons. We plan to investigate these effects in the future using the pinhole algorithm [Elhatisari:2017eno], see also the recent work in [Ren:2025vpe]. However, pinhole calculations of $A$-body density correlations do not have an immediate analog for other nuclear many-body methods. It is therefore valuable that a significant amount of information about nuclear forces and quantum correlations can be deduced from the simple correlation measurements introduced here and can be expressed in a model-independent language. The correlation studies presented here can be readily adopted by other groups using different nuclear many-body methods.